# Reason-to-Rank: Learning to Rank through Reasoning-Based Knowledge Distillation

### v2

@@
Abstract
Reranking documents based on their relevance to a given query is a critical task in information retrieval. Traditional reranking methods often lack transparency and rely on proprietary models, hindering reproducibility and interpretability. We propose Reason-to-Rank (R2R), a novel open-source reranking approach that enhances transparency by generating two types of reasoning: direct relevance reasoning, which explains how a document addresses the query, and comparison reasoning, which justifies the relevance of one document over another. We leverage large language models (LLMs) as teacher models to generate these explanations and distill this knowledge into smaller, openly available student models. Our student models are trained to generate meaningful reasoning and rerank documents, achieving competitive performance across multiple datasets, including MSMARCO and BRIGHT. Experiments demonstrate that R2R not only improves reranking accuracy but also provides valuable insights into the decision-making process. By offering a structured and interpretable solution with openly accessible resources, R2R aims to bridge the gap between effectiveness and transparency in information retrieval, fostering reproducibility and further research in the field. [^1] Our code and data are available here https://anonymous.4open.science/r/Distillation-Is-All-You-Need-C98D/README.md
Reason-to-Rank: Learning to Rank through Reasoning-Based Knowledge Distillation
Yuelyu Ji1, Zhuochun Li1, Rui Meng2, Daqing He1 1School of Computing and Information, University of Pittsburgh, 2Salesforce Research {yuj49,zhl163, dah44}@pitt.edu, ruimeng@salesforce.com
1 Introduction
The ability to effectively rank and rerank documents is critical in information retrieval (IR), where the quality of search results significantly impacts user satisfaction Pradeep et al. (2023b); Sun et al. (2023). Traditional reranking approaches often improve initial rankings provided by retrieval systems but pay limited attention to the reasoning behind ranking decisions. Recent studies have revealed that incorporating reasoning into the reranking process can enhance performance and interpretability Pradeep et al. (2023b); Liu et al. (2024); Yu et al. (2024c); Sun et al. (2023). By offering explanations that clarify why certain documents are prioritized over others, these methods help users understand and trust the search results.
However, most existing reranking models do not distinguish between two key types of reasoning: direct relevance reasoning, which explains how a document matches a query, and comparative reasoning, which justifies why one document ranks higher than another. This limitation can reduce ranking effectiveness, particularly in contexts where a subtlety understanding of document relevance is crucial.
@@ 1 Introduction
Figure 1: Illustration of the two types of reasoning: direct relevance reasoning provides explicit answers to the query, while comparison reasoning evaluates the relative relevance between documents. The LLM generates these explanations to enhance the interpretability of the ranking process.
In this paper, we propose Reason-to-Rank (R2R), a reranking framework that leverages both explicit and comparative reasoning to provide more detailed and transparent document rankings. Our approach based on knowledge distillation, where a large, powerful teacher model (an LLM) generates high-quality explanations and rankings. Distilling these explanations into smaller, more computationally efficient student model that perform the reranking tasks. Our contributions are threefold:
  • Introduce a reranking approach that separates direct relevance and comparative reasoning, enhancing the clarity of document rankings.
  • Demonstrate that student models can learn from LLM-generated data, achieving high reranking performance while remaining computationally efficient.
  • Provide an open-source implementation to facilitate further research and applications in information retrieval.
The remainder of this paper is organize as follows. §2 reviews related work. §3 presents the Reason-to-Rank (R2R) framework. Sections 4 not only present the quantitative performance metrics but also provide a thorough discussion on the implications of the results, analyzing how the reasoning strategies impact the ranking performance and offering insights into the model’s interpretability. Finally, §6 concludes the paper and discusses future work.
2 Related Work
The evolution of reranking techniques has significantly advanced with the rise of neural networks, particularly transformers Vaswani (2017); Yu et al. (2024b, a); Zheng et al. (2024); Zhang et al. (2024); Shen et al. (2024); Fan and Tao (2024); Fan et al. (2024). Early models like monoBERT Nogueira et al. (2019) utilized BERT Devlin (2018) in a cross-encoder setup, effectively capturing contextual dependencies between queries and documents. Further models such as monoT5 Nogueira et al. (2020) and InPars Bonifacio et al. (2022) continued to push reranking performance boundaries, especially on datasets like MS MARCO Bajaj et al. (2016). However, these models rely heavily on large-scale fine-tuning and task-specific data, limiting their scalability and adaptability to new domains or tasks.
@@ 2 Related Work
Figure 2: Overview of direct relevance and comparison reasoning prompts for document ranking. Pairwise reasoning explains how a document matches a query, while comparison reasoning evaluates the relative relevance between documents.
@@
Figure 3: Overview of the Reason-to-Rank framework. The teacher model generates direct relevance and comparative reasoning, which is used to train student models capable of reproducing the reranked order and generating explanations.
Early ranking systems primarily focused on direct relevance methods, where each document’s relevance was judged in isolation Liu et al. (2009). However, pairwise approaches like duoBERT Pradeep et al. (2021) were introduced to compare pairs of documents, determining which document was more relevant. While this improved the relative ranking of documents, it still missed the larger picture of the entire ranked list. To address this, listwise models such as RankGPT Sun et al. (2023) and its distilled versions like RankVicuna Pradeep et al. (2023a) considered the entire ranked list during training, allowing for more holistic improvements in ranking accuracy. However, even these methods have limitations when it comes to explaining why a particular document ranks higher than another.
The Learning to Rank framework introduced several foundational approaches in information retrieval, including pointwise, pairwise, and listwise methods, which form the basis of modern neural ranking models Liu et al. (2009). Direct relevance approaches focus on predicting the relevance score of individual documents in isolation, while pairwise methods aim to indicate the relative ranking between document pairs. On the other hand, listwise approaches focus on directly optimizing the entire ranking list, offering better overall accuracy by considering the global structure of the ranking. Despite these advances, most early methods lacked clarity that modern applications demand.
Recent work has also highlighted the importance of natural language explanations in improving clarity and trust in reranking decisions. Models such as Explain-then-Rank Zhuang et al. (2023) have focused on generating explanations that describe how a document matches a query. However, these approaches often only address direct relevance reasoning (i.e., document-query matching) and do not account for comparative rationale (i.e., why one document ranks higher). Our work builds on these efforts by incorporating both types of reasoning, enhancing the overall clarity and coherence of the explanations. This dual focus provides more precise insights into the ranking process and improves the model’s ability to deliver accurate and logically consistent rankings Doshi-Velez and Kim (2017); Lipton (2018); Mo et al. (2024); Yu et al. (2024b, a).
Knowledge distillation has become an essential technique in training smaller, more efficient student models from larger teacher models Hinton et al. (2015) to further reduce the computational demands of large reranking models. Knowledge distillation has been particularly effective in reranking tasks, as shown by models like RankVicuna Pradeep et al. (2023a), which distilled the reranking capabilities of RankGPT into a smaller, more efficient model without significant loss of performance. Our approach builds on this by applying distillation to both the reasoning and ranking components, ensuring that our student models not only maintain high performance but also offer explainability in their reranking decisions Qin et al. (2023); Dai et al. (2022); Xu et al. (2024b, a).
By integrating these advances, our R2R framework improves existing methods by offering a dual reasoning strategy - direct relevance and comparative reasoning - within a computationally efficient model. Our approach not only enhances ranking accuracy but also provides clear and easy-to-understand explanations for the ranking decisions, addressing fundamental limitations in performance seen in prior work.
3 Method
3.1 Teacher Model with Dual Reasoning
@@
Given a query $q$ and a set of candidate documents $\{d_{1},d_{2},\dots,d_{n}\}$, the teacher model processes these inputs to generate two types of reasoning and rerank order. The teacher model can generate high-quality reasoning that combines both direct relevance and comparative aspects, and the examples can be seen in Figure 2: Direct relevance Reasoning: The model explains how each document directly addresses the query, focusing on the relevance and specificity of the content. Comparative Reasoning: The model assesses the relative relevance between documents, explaining why one document should be ranked higher based on content.
3.2 Student Model with Knowledge Distillation
@@
Deploying large language models (LLMs) for reranking is often impractical due to their high computational cost. To address this, we design the student model to replicate the teacher’s performance while being more efficient for real-world deployment. The innovation lies in using knowledge distillation: the student learns from the teacher’s reasoning and ranking outputs, retaining accuracy and interpretability with significantly lower computational demands. This approach ensures the student can perform complex reranking tasks efficiently, making it ideal for large-scale use.
3.2.1 Model Overview
@@ 3.2.2 Training Objective
Next, the listwise loss optimizes the overall ranking of the document list. It uses Kullback-Leibler (KL) divergence Hershey and Olsen (2007) to measure the difference between the predicted ranking distribution and the target ranking distribution:
@@
where $P(\mathbf{s})=\text{softmax}(\mathbf{s}_{i})$ is the predicted ranking distribution based on the model’s output, and $Q(\mathbf{z})=\text{softmax}(\mathbf{z}_{i})$ is the target ranking distribution based on true labels.
Finally, the generation loss optimizes the quality of the textual reasoning generated by the model. It is based on the cross-entropy loss between the generated reasoning and the ground-truth reasoning:
@@
DeBERTa Sun et al. (2023) 68.5 64.2 47.0 73.4 32.1 50.2 33.7 49.2 45.4
MonoT5 Nogueira et al. (2020) 74.5 70.4 48.3 80.0 34.1 46.0 35.2
RankVicuna Pradeep et al. (2023a) 71.1 68.7 48.3 67.1 48.7 38.5 55.7 35.3
RankZephyr Pradeep et al. (2023b) 72.2 70.5 51.3 85.1 36.5 53.3 38.9 60.7 35.5
APEER Jin et al. (2024) 74.6 72.3 51.1 83.9 35.3 52.1 33.4 56.0 46.1
@@
Table 1: NDCG@5 performance (in percentage) for student models and baseline comparisons across multiple datasets. ’w/o Reasoning’ models do not use reasoning prompts and do not generate explanations; they are optimized using the specified ranking loss functions. ’w/ direct relevance Reasoning’ and ’w/ comparison Reasoning’ models use reasoning prompts and include reasoning loss for generating explanations. ’’ indicates statistically significant improvement over the baseline DeBERTa model (p $<$ 0.05). Bold indicates the best performance among all models for each dataset. Entries marked with ’–’ indicate that results are not available for that dataset.
4 Experiments
@@
4.1 Datasets and Evaluation Metrics
In addition to MSMARCO Bajaj et al. (2016) and BEIR Thakur et al. (2021), we utilize the BRIGHT Su et al. (2024) to further evaluate our model’s ability to handle reasoning-intensive retrieval tasks.

### v4

# Reason-to-Rank: Distilling Direct and Comparative Reasoning from Large Language Models for Document Reranking

@@
Abstract.
Reranking documents in information retrieval often relies on black-box models that improve effectiveness but lack explainability. We introduce Reason-to-Rank (R2R), a novel framework that separates direct relevance reasoning from comparison reasoning to provide both direct and comparitive explanations. We first prompt a large language model to produce comprehensive rationales and a ranking order; then we distill both the ranking decisions and textual explanations into a smaller, open-source student model. Our approach not only improves retrieval performance, as demonstrated in MSMARCO, BEIR, and BRIGHT, but also provides interpretable justifications for why one document outranks another. We report NDCG@5 (and NDCG@10) for direct comparisons with prior work, and show that the distilled student model achieves competitive results while significantly reducing computational overhead. By unifying direct and comparative reasoning in a single pipeline, R2R bridges the gap between transparency and effectiveness in modern reranking systems.
Large Language Models; Information Retrieval; Re-ranking
journalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730070isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Learning to rank
1. Introduction
@@
Figure 1. Illustration of the two types of reasoning: direct relevance reasoning provides explicit answers to the query, while comparison reasoning evaluates the relative relevance between documents. The LLM generates these explanations to enhance the interpretability of the ranking process.
With the explosive growth of online information, modern search engines and recommendation systems heavily rely on reranking to refine initial retrieval results and present users with the most relevant content (Sun et al., 2023; Pradeep et al., 2023b). Despite significant progress in neural reranking, particularly with large language models (LLMs) such as GPT-4, contemporary solutions often lack transparency in how they differentiate between highly similar candidate documents. Users see a final ranking list, yet remain uncertain about why how specific items are prioritized, which can reduce trust and hinder effective system debugging (Nogueira et al., 2020; Liu et al., 2024).
Explainable reranking has attracted increasing attention, with the aim of combining ranking decisions with human-readable justifications. Recent studies (Yu et al., 2024; Zhuang et al., 2023) show that providing textual explanations can improve user understanding and confidence, and can also help system designers pinpoint potential errors in the ranking pipeline. However, most prior approaches focus on a single dimension of explanation, direct relevance reasoning, where each document is described in isolation: “This document is relevant to the query because of…”. While useful, such direct rationales do not address the relative importance of documents, a critical aspect when multiple items appear similarly relevant yet differ in subtle ways.
We argue that comparison reasoning is the missing piece in many existing explainable reranking systems. By comparing pairs (or subsets) of documents—explaining, for instance, why Document A is placed above Document B—the model can highlight unique advantages or deficiencies, enabling more fine-grained control over the final ranked list (Pradeep et al., 2021; Sun et al., 2023). Without such pairwise or listwise clarity, users and IR practitioners must rely on guesswork when interpreting top results, limiting the overall efficacy of explanation-based retrieval solutions.
In this paper, we propose Reason-to-Rank (R2R), a novel framework that unifies direct relevance with comparison-based reasoning. R2R leverages a powerful teacher LLM (e.g., GPT-4) to produce both a global ranking and two kinds of textual justifications: (1) a direct explanation for each candidate, and (2) a comparative explanation indicating which document outranks another and why. The framework is shown in Figure 1.We then employ a knowledge distillation procedure to train a smaller student model to replicate the teacher’s ranking decisions and explanations. From an educational perspective, research shows that learners benefit more when they are taught not just answers but the reasoning behind them (Raza, 2020). Inspired by this, our approach transfers both the what and the why, helping the student model generalize better and generate its own explanations with improved clarity. This setup offers two major advantages: (i) The final deployed system preserves interpretability without incurring the high computational cost of querying a large LLM in production, and (ii) By explicitly encoding comparative rationales, R2R achieves more subtlety distinctions among the top candidates than direct relevance only methods.
We perform extensive experiments on multiple benchmarks, including MSMARCO (Bajaj et al., 2016), BEIR (Thakur et al., 2021), and BRIGHT (Su et al., 2024), covering both general search queries and reasoning-intensive tasks. The results show that R2R outperforms various baselines (e.g., monoT5 (Nogueira et al., 2020), RankVicuna (Pradeep et al., 2023a)) in terms of NDCG metrics, while also providing more interpretable output. A comprehensive ablation study further validates the importance of integrating both direct relevance and comparison reasoning, as omitting either dimension leads to decreased performance and reduced clarity. In addition, we provide qualitative examples illustrating how comparison explanations help reveal subtle but important differences among retrieved documents.
In summary, our primary contributions are:
  • We introduce a novel reranking framework, R2R, that produces both direct relevance and comparison-based explanations, offering a more transparent and fine-grained view of document ranking.
  • We propose a teacher–student knowledge distillation pipeline, where a large LLM generates enriched multiperspective rationales, subsequently transferred to a smaller, more efficient student model for practical deployment.
  • We conduct extensive evaluations on general and reasoning-focused datasets, demonstrating that R2R achieves not only competitive or superior reranking performance but also enhanced interpretability through dual-type reasoning.
@@
Figure 2. Overview of direct relevance and comparison reasoning prompts for document ranking. Direct relevance reasoning explains how a document matches a query, while comparison reasoning evaluates the relative relevance between documents.
2. Related Work
2.1. Neural Reranking and Explanation
Neural reranking has been extensively studied to refine candidate lists obtained from first-stage retrieval (e.g., BM25 or dense retrievers) (Nogueira et al., 2019, 2020). Early efforts often used cross-encoder architectures (e.g., BERT or T5) to capture fine-grained query–document interactions. Although these models achieve strong effectiveness, they typically present limited insights regarding the reasoning behind their decisions (Liu et al., 2009).
Recent work has thus explored generating explanations alongside the ranking (Zhuang et al., 2023; Yu et al., 2024). For instance, Explain-then-Rank (Zhuang et al., 2023) adds textual justifications on how a document addresses the query. However, these methods tend to focus on direct relevance (i.e., how well a single document matches the query) rather than explicitly modeling comparative relevance among documents, which may be crucial in distinguishing subtle differences at the top of the ranked list (Doshi-Velez and Kim, 2017; Lipton, 2018).
2.2. Direct vs. Comparison Reasoning
Classical Learning to Rank (LTR) frameworks (Liu et al., 2009) have shown that pointwise, pairwise, and listwise strategies each possess unique advantages. Pairwise approaches like duoBERT (Pradeep et al., 2021) are explicitly designed to compare two documents and decide which one is more relevant, whereas pointwise methods focus on rating each document in isolation. More recently, listwise approaches (e.g., RankGPT (Sun et al., 2023) or RankVicuna (Pradeep et al., 2023a)) consider the entire ranked list, often achieving stronger empirical results. Nonetheless, even these advanced listwise techniques do not always clearly explain why one document outranks another. Our proposed R2R differs by explicitly producing both direct relevance and comparative rationales within a unified framework.
2.3. Knowledge Distillation in Reranking
Recent work has addressed the dual goals of ranking effectiveness and transparency. On the explanation front, EXS (Singh and Anand, 2019) offers local, model-agnostic rationale extraction for pointwise rankers, whereas RankingSHAP (Heuss et al., 2024) applies Shapley-value analysis across entire ranked lists. Both methods clarify why documents are preferred, but do not compress these insights into smaller, more efficient models.
Meanwhile, methods like One-Shot Labeling (MacAvaney and Soldaini, 2023) and Few-Shot Prompting for Pairwise Ranking (Sinhababu et al., 2024) demonstrate that even sparse or minimal relevance data can significantly boost retrieval. Although they refine ranking accuracy via limited supervision, they typically do not target explanation generation or model compression.
Overall, our approach is complementary: we incorporate both direct and comparative reasoning, and distill these insights into an open-source student model that can both rank and explain. Extending Reason-to-Rank to incorporate advanced explanation techniques such as Shapley-based analysis is a promising avenue for future work.
Figure 3. Basic prompt and return type.
@@ 2.3. Knowledge Distillation in Reranking
Figure 4. Overview of the Reason-to-Rank framework. The teacher model generates direct relevance and comparative reasoning, which is used to train student models capable of reproducing the reranked order and generating explanations.
3. Method
Initial Retrieval with BM25.
Following common practice in neural re-ranking (Nogueira et al., 2019; Sun et al., 2023), we first retrieve the top-100 candidate documents per query using BM25. This initial retrieval step serves as the candidate pool from which our re-ranking pipeline will operate. We choose 100 to balance recall (ensuring that most relevant documents are included) and computational overhead. [^1] Our code and data are available here https://github.com/JoyDajunSpaceCraft/Reason-to-Rank.git
3.1. Teacher Model with Dual Reasoning
@@
Given a query $q$ and a set of candidate documents $\{d_{1},d_{2},\dots,d_{n}\}$, the teacher model processes these inputs to generate two types of reasoning and rerank order. The teacher model can generate high-quality reasoning that combines both direct relevance and comparative aspects, and the examples can be seen in Figure 2: Direct relevance Reasoning: The model explains how each document directly addresses the query, focusing on the relevance and specificity of the content. Comparative Reasoning: The model assesses the relative relevance between documents, explaining why one document should be ranked higher based on content. We also have the basic prompt and the return type in Figure 3. Our teacher model (e.g., GPT-4) processes the top-100 candidates and produces two outputs:
  1. (1)
    A ranking order (i.e., a global ordering of the 100 documents), which corresponds to how the teacher prioritizes each document’s relevance to the query.
  2. (2)
    A textual explanation for each document, covering both direct relevance (i.e., how the document addresses the query) and comparison reasoning (i.e., why one document is more relevant than another).
While the teacher internally may compute or utilize relevance scores, we unify these into a final ordering to avoid ambiguity. In other words, rather than exposing raw numeric scores, the teacher model outputs a definitive rank list (e.g., from most relevant to least relevant) alongside the corresponding rationales.
3.2. Student Model with Knowledge Distillation
@@
Deploying large language models (LLMs) for reranking is often impractical due to their high computational cost. To address this, we design the student model to replicate the teacher’s performance while being more efficient for real-world deployment. We obtain numeric ranking scores for supervision by prompting the teacher to produce a ranked list and converting the document positions into scores via normalized ranks. The innovation lies in using knowledge distillation: the student learns from the teacher’s reasoning and ranking outputs, retaining accuracy and interpretability with significantly lower computational demands. This approach ensures the student can perform complex reranking tasks efficiently, making it ideal for large-scale use.
3.2.1. Model Overview
@@ 3.2.2. Training Objective
Next, the listwise loss optimizes the overall ranking of the document list. It uses Kullback-Leibler (KL) divergence (Hershey and Olsen, 2007) to measure the difference between the predicted ranking distribution and the target ranking distribution:
@@
where $P(\mathbf{s})=\text{softmax}(\mathbf{s}_{i})$ is the predicted ranking distribution based on the model’s output, and $Q(\mathbf{z})=\text{softmax}(\mathbf{z}_{i})$ is the target ranking distribution based on true labels. We treated the pairwise loss and listwise loss as rank loss. Finally, the generation loss optimizes the quality of the textual reasoning generated by the model. It is based on the cross-entropy loss between the generated reasoning and the ground-truth reasoning:
@@
Table 1. NDCG@5 performance (in percentage) for different models with various reasoning prompts across multiple datasets. Bold numbers indicate the best performance within each group of models. indicates statistically significant improvement over the base model without reasoning prompts (p $<$ 0.05). indicates statistically significant improvement over both the base model and the single reasoning prompt variants (p $<$ 0.05).DL19 and DL20 (Craswell et al., 2020) are test datasets for MSMARCO. The MonoT5 has no result for News, and NFCorpus and RankVicuna has no results for News in their paper.
BM25 50.6 48.0 41.0 59.5 44.2 39.5 30.8 40.7 31.8
Rank GPT-4 Model
Rank GPT-4 (w/o reasoning) 75.6 70.6 54.0 85.5 38.5 57.6 36.6 57.6 47.1
+ direct relevance reasoning 75.3 69.9 52.9 83.3 38.9 55.9 36.1 52.9 47.0
+ comparison reasoning 76.3 71.1 53.4 84.6 38.2 56.5 36.5 56.5 48.2
Claude Model
Claude (w/o reasoning) 72.0 68.8 50.8 82.8 36.1 54.4 36.4 50.8 42.5
+ direct relevance reasoning 70.6 68.2 50.3 82.2 35.9 55.4 35.6 52.4 42.5
+ comparison reasoning 71.3 69.8 51.1 83.3 36.4 55.1 36.7 51.5 43.2
Gemini Model
Gemini (w/o reasoning) 71.1 68.5 50.1 83.0 35.6 53.7 35.5 50.1 42.9
+ direct relevance reasoning 69.1 67.0 51.1 81.6 34.8 52.9 36.9 51.3 43.9
+ comparison reasoning 70.2 68.2 50.2 82.4 35.9 53.4 35.4 51.1 42.9
Table 2. NDCG@5 performance (in percentage) for student models and baseline comparisons across multiple datasets. ’w/o Reasoning’ models do not use reasoning prompts and do not generate explanations; they are optimized using the specified ranking loss functions. ’w/ direct relevance Reasoning’ and ’w/ comparison Reasoning’ models use reasoning prompts and include reasoning loss for generating explanations. ’’ indicates statistically significant improvement over the baseline DeBERTa model (p $<$ 0.05). Bold indicates the best performance among all models for each dataset. Entries marked with ’–’ indicate that results are not available for that dataset.
@@
DeBERTa (Sun et al., 2023) 68.5 64.2 47.0 73.4 32.1 50.2 33.7 49.2 45.4
MonoT5 (Nogueira et al., 2020) 74.5 70.4 48.3 80.0 34.1 46.0 35.2
RankVicuna (Pradeep et al., 2023a) 71.1 68.7 48.3 67.1 48.7 38.5 55.7 35.3
RankZephyr (Pradeep et al., 2023b) 72.2 70.5 51.3 85.1 36.5 53.3 38.9 60.7 35.5
APEER (Jin et al., 2024) 74.6 72.3 51.1 83.9 35.3 52.1 33.4 56.0 46.1
@@
4. Experiments
@@
4.1. Datasets and Evaluation Metrics
In addition to MSMARCO (Bajaj et al., 2016) and BEIR (Thakur et al., 2021), we utilize the BRIGHT (Su et al., 2024) to further evaluate our model’s ability to handle reasoning-intensive retrieval tasks.